The postion x of a particle moving along x-axis varies with time t as x = Asin(αt) where a and α are positive constants . the acceleration a of the partcle varies with position (x) as

Grade:12

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
\\ \\x=Asin(\alpha t) \\ {dx^2 \over dt^2}=A\alpha ^2(-sin(\alpha t)) \\a=A\alpha ^2(-x) \\
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty